OpenStudy (logan13):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (logan13):

more than 1

OpenStudy (logan13):

@paki

OpenStudy (anonymous):

I think you may need to review what closed under an operation means. 1 + 1 = 2, which isn't in the set. Therefore we can get outside of the set by adding elements. When we can do this, it isn't closed under addition. 1 - (-1) = ? 1 / 0 = ? As it turns out, of those four operations, the set is only closed under multiplication

OpenStudy (anonymous):

does that help any?

OpenStudy (logan13):

@lilshane